A fidget spinner of radius 0.05 m has a moment of inertia kg m2

and


is rotating at 8.4 rad s-1


. A large fly whose mass is 0.3 kg lands onto the edge of

the rotating spinner. Calculate the angular speed of the system after the fly falling

onto the spinner.


1
Expert's answer
2021-10-21T16:33:46-0400

Given:

"r=0.05\:\rm m"

"I_0=1.25*10^{-4}\:\rm kg\cdot m^2"

"\omega_1=8.4\:\rm rad/s"

"m=0.3\:\rm kg"


The law of conservation of angular momentum

"I_0\omega_1=(I_0+mr^2)\omega_2"

Hence


"\omega_2=I_0\omega_1/(I_0+mr^2)\\
=1.25*10^{-4}*8.4/(1.25*10^{-4}+0.3*0.05^2)\\
=1.2\:\rm rad/s"